We are starting multistep equations. I am lost when the equation has the variable as part of a fraction on both sides of the equation. For instance:
b/5 + 8 = b/2 + 10

also:
t + 2/3 + 5t/6 = 4t - 5/12
Thank you!
Cassie

Answer by stanbon(75887)

b/5 + 8 = b/2 + 10
Find a "least-common-multiple" of the denominators and multiply thru by it
to get rid of of the fractions.
Your LCM is 10, so:
2b+80=5b+100
Get the variable terms on one side and the numbers on the other:
3b=-20
b=-20/3
----------------
also:
t + 2/3 + 5t/6 = 4t - 5/12
The LCM is 12 so multiply thru by it:
12t+8+10t=48t-5
26t=13
t=13/26
t=1/2
--------------
Cheers,
Stan H.